# Exploring 3D Activity Reasoning and Planning: From Implicit Human Intentions to Route-Aware Planning

### v1

@@ 4.2 Overall Framework
4.3 Progressive Plan Generation
Contextual information is essential for ensuring consistency across steps in our activity planning task. To this end, we propose a progressive plan generation approach that explicitly enforces constraints among different steps by leveraging previously generated plans. As illustrated in Algorithm 1, when generating the first step, we provide the human’s implicit instruction $X_{inst}$ as input to the LLM. For subsequent steps, we prompt the LLM with an input historical step guidance template $X_{h}$, which incorporates previously generated plans and explicitly instructs the LLM to follow them. Specifically:
4.4 Dynamic Graph Modulation
We propose a dynamic graph modulation mechanism that adaptively adjusts the importance of nodes and edges in the generated scene graph, thereby enhancing the awareness of relevant objects in the generated action plan. Specifically, the generated scene graph $G$ consists of nodes $N$ (representing objects) and edges $E$ (representing spatial relations between two objects). When the generated one-step plan $\hat{Y}_{plan}^{s}$ includes the embedding of $i$-th object $N_{i}$, we emphasize $N_{i}$, its $K$ neighbors and their spatial relations by applying a higher weight $w_{h}$ to the embeddings, as formulated below:
$\{N_{i},N_{ij},E_{ij}\}*w_{h},j\in\{1,\cdots,K\}$ (1)
where node $N_{ij}$ denotes the $j$-th neighbor of $N_{i}$, and edge $E_{ij}$ represents the spatial relation between $N_{ij}$ and $N_{i}$. This mechanism can effectively coordinate with the proposed Progressive Plan Generation approach in the iterative process, as illustrated in Algorithm 1.
By focusing on the most relevant objects, their neighbors, and corresponding spatial relations, this dynamic graph modulation mechanism improves the awareness of critical objects and their spatial relations which is beneficial for route planning within our task.
4.5 Training Objective
$L_{plan}=CE(Y_{plan},\hat{Y}_{plan}),$ (2)
Figure 5: Activity reasoning and planning visualization over the ReasonPlan3D val set. Each example shows an implicit human instruction, the input point clouds of the 3D scene, the 3D segmentation of the scene, and the predictions from 3DGraphLLM and the proposed SHARP. Best viewed in color and zoom-in.
5 Experiment
5.1 Experimental Settings
Evaluation Metrics.
We adopt BLEU [37], ROUGE [32], METEOR [3], and CIDEr [49] to evaluate the quality of generated plans for the reasoned activities. In addition, we adopt BLEU-4, CIDEr, and METEOR to measure the performance of activity reasoning.
Implementation Details.
We train our framework for $11k$ iterations with a batch size of $2$. We employ the AdamW [33] optimizer with an initial learning rate of $2e^{-5}$, a weight decay of $0.02$, and a cosine scheduler with $1.1k$ warm-up iterations. We adopt LLaMA-3-8B-Instruct [1] as our multimodal LLM backbone and apply LoRA [19] with a rank of 16 for efficient fine-tuning. The weight $w_{h}$ used in the proposed Dynamic Graph Modulation is set as $2.0$.
@@ 5.2 Benchmarking with Existing Methods
Table 2 presents quantitative experiments on the ReasonPlan3D validation set. We can observe that the proposed SHARP achieves superior performance across all evaluation metrics, demonstrating its superiority in activity reasoning, task planning, and inter-step route planning. The superior performance is largely attributed to two key factors. First, the proposed Progressive Plan Generation enforces the model to maintain contextual consistency throughout the plan generation process by referring to the plans generated in previous steps. Second, the proposed Dynamic Graph Modulation guides the model toward critical objects and their spatial relations, enhancing 3D scene understanding and boosting route planning accuracy.
@@
Table 3: Benchmarking on the ReasonPlan3D validation set for measuring the performance of activity reasoning with evaluation metrics BLEU-4, CIDEr, and METEOR. Best in bold, second underlined.
In addition, Table 3 presents the activity reasoning performance of SHARP and several state-of-the-art methods over the validation set of ReasonPlan3D. Our approach achieves superior performance across all metrics. Specifically, 3D-LLM [18], LL3DA [6], and Chat-Scene [21] do not employ a scene graph to explicitly guide the model’s awareness of critical objects, leading to suboptimal performance when dealing with the activity reasoning task where comprehending object relations plays an essential role. 3DGraphLLM [56] employs only scene graphs without scene-level embeddings of the global picture, leading to inferior 3D scene understanding and weaker performance in the activity reasoning task.
@@
Index $w_{h}$ BLEU-1 BLEU-2 BLEU-3 BLEU-4 CIDEr METEOR ROUGE
@@
Table 5: Ablation studies on the weight $w_{h}$ as used in the proposed Dynamic Graph Modulation. The best results are in bold.
Qualitative Benchmarking.
@@ 5.3 Ablation Study
We conduct ablation studies over the validation set of ReasonPlan3D to evaluate the effectiveness of our designs, including Progressive Plan Generation and Dynamic Graph Modulation as presented in the SHARP framework. In addition, we examine the effect of the weight $w_{h}$ as used in the Dynamic Graph Modulation.
@@
Effectiveness of $w_{h}$.
As defined in Equation 1, the weight $w_{h}$ in the Dynamic Graph Modulation is introduced to highlight critical objects and their spatial relations relevant to the reasoned activity. We examine the effectiveness of $w_{h}$ in Table 5. It can be observed that critical objects and their spatial relations are not well captured when $w_{h}$ is set as $1.0$, leading to degraded performance. Increasing $w_{h}$ to 1.5 improves consistently, suggesting that increasing emphasis on critical objects and their spatial relations can boost the model’s planning ability. Further increasing $w_{h}$ to 2.0 achieves the best performance, confirming that guiding the model toward these crucial cues effectively enhances the model’s planning ability. However, increasing $w_{h}$ to 3.0 overemphasizes these objects and their spatial relations, potentially leading to an imbalance among objects in the 3D scene that degrades the performance.

### v2

@@ 5.2 Benchmarking with Existing Methods
Table 2 presents quantitative experiments on the ReasonPlan3D validation set, where all compared methods are trained and evaluated on the same ReasonPlan3D benchmark for fairness. We can observe that the proposed SHARP achieves superior performance across all evaluation metrics, demonstrating its superiority in activity reasoning, task planning, and inter-step route planning. The superior performance is largely attributed to two key factors. First, the proposed Progressive Plan Generation enforces the model to maintain contextual consistency throughout the plan generation process by referring to the plans generated in previous steps. Second, the proposed Dynamic Graph Modulation guides the model toward critical objects and their spatial relations, enhancing 3D scene understanding and boosting route planning accuracy.
In addition, Table 3 presents the activity reasoning performance of SHARP and several state-of-the-art methods over the validation set of ReasonPlan3D. Our approach achieves superior performance across all metrics. Specifically, 3D-LLM [18], LL3DA [5], and Chat-Scene [20] do not employ a scene graph to explicitly guide the model’s awareness of critical objects, leading to suboptimal performance when dealing with the activity reasoning task where comprehending object relations plays an essential role. 3DGraphLLM [56] employs only scene graphs without scene-level embeddings of the global picture, leading to inferior 3D scene understanding and weaker performance for activity reasoning.
Figure 5: Activity reasoning and planning visualization over the ReasonPlan3D val set. Each example shows an implicit human instruction, the input point clouds of the 3D scene, the 3D segmentation of the scene, and the predictions from 3DGraphLLM and SHARP. Best viewed in color and zoom-in.
Table 3: Benchmarking on the ReasonPlan3D validation set for measuring activity reasoning performance with evaluation metrics BLEU-4, CIDEr, and METEOR. Best in bold, second underlined.
@@
Qualitative Benchmarking.
@@
Table 5: Ablation studies on the weight $w_{l}$ as used in the proposed Dynamic Graph Modulation. The best results are in bold.
Index $w_{l}$ BLEU-1 BLEU-2 BLEU-3 BLEU-4 CIDEr METEOR ROUGE
@@ 5.3 Ablation Study
We conduct ablation studies over the validation set of ReasonPlan3D to evaluate the effectiveness of our designs, including Progressive Plan Generation and Dynamic Graph Modulation as presented in the SHARP framework. In addition, we examine the effect of the weight $w_{l}$ as used in the Dynamic Graph Modulation.
@@
Effectiveness of $w_{l}$.
As defined in Equation 1, the weight $w_{l}$ in the Dynamic Graph Modulation is introduced to highlight critical objects and their spatial relations relevant to the reasoned activity. We examine the effectiveness of $w_{l}$ in Table 5. It can be observed that critical objects and their spatial relations are not well captured when $w_{l}$ is set as $1.0$, leading to degraded performance. Increasing $w_{l}$ to 1.5 consistently improves performance, suggesting that placing greater emphasis on critical objects and their spatial relations can boost the model’s ability. Further increasing $w_{l}$ to 2.0 achieves the best performance, confirming that guiding the model toward these crucial cues enhances the model’s planning ability. However, increasing $w_{l}$ to 3.0 overemphasizes these objects and their spatial relations, leading to an imbalance among objects in the 3D scene that degrades the performance.